# Beyond the Main Sequence: Binary Evolution Pathways to Blue Stragglers in the Gaia Era

### v1

@@ Abstract
Context. The study of blue straggler stars (BSS) provides insight into the mechanisms of stellar mass exchange during binary stellar evolution and the complex gravitational interactions within dense stellar systems. In combination, they enhance our understanding of the possible life cycles of stars and the evolutionary pathways of star clusters.
Aims. We study the populations of BSSs in 41 globular clusters (GCs) and 42 open clusters (OCs) adopting photometry, proper motion, and parallax from the Gaia Data Release 3 (DR3), confirming their cluster membership. We find a total of 4399 BSSs: all GCs show the presence of BSSs (3965 or of the sample), whereas only 42 out of 129 studied OCs show the presence of BSSs (434 or of the sample). Clusters younger than Myr do not host any BSSs.
Methods. We derive their astrophysical parameters such as effective temperature, surface gravity, and mass based on color-temperature relations, isochrone models and Gaia DR3 spectroscopy (if available). We find values for and ; and an average mass of and for GCs and OCs, respectively. We finally compute the difference of the BSS mass and the main sequence turn-off (MSTO) mass of its respective cluster, normalized by the MSTO mass, for every identified BSS.
Results. Based on this parameter and BSSs ages derived from isochrone models, we find: i) GC BSSs that are most likely to be formed through collisions show a “boost” for ages Gyr, in agreement with ages for core-collapse events in GCs reported in previous studies; ii) a double sequence for GC BSSs that could indicate the hint of a pre/post-merger or close-binary scenario.
@@
blue stragglers – globular clusters: general – open clusters and associations: general1 Introduction
1.1 Overview and Formation Mechanisms
@@
Blue Stragglers Stars (BSSs) are stars that appear to be younger and more massive than the bulk of other stars in their host stellar population. They are characterized by their location in the color-magnitude diagram (CMD), where they appear hotter and bluer than their main-sequence (MS) neighbours.
@@
BSSs were first identified by Sandage (1953) through his observations of the globular cluster (GC) Messier 3 (M3) and some years later they were reported in the open cluster (OC) NGC 7786 by Burbidge & Sandage (1958). Ever since, their presence has been noticed in distinct structures: as field stars (Preston & Sneden, 2000; Jofré et al., 2016), in OB associations (Mathys, 1987), the Galactic bulge (Clarkson et al., 2011) and in dwarf galaxies (Da Costa, 1984; Mapelli et al., 2009).
@@
BSSs put the theory of stellar evolution in “check”: i.e. the more massive a star is, the faster it evolves. For this reason, these stars should have evolved off the MS branch a long time ago; but we still find them populating this branch as an apparent extension of the MS in a CMD. Three scenarios have been proposed to explain their formation: i) Mass-transfer (MT) in binary systems (McCrea, 1964), where a star is receiving material from its companion, extending its life; ii) Stars formed through collisions/mergers (Hills & Day, 1976; Lombardi et al., 2002); and iii) Hierarchical triple systems: magnetized winds or stellar evolution cause the loss of angular momentum in a multiple system, which leads to collisions (or MT) between the stars that conform the system (Perets & Fabrycky, 2009; Naoz & Fabrycky, 2014). It has been shown that the BSS properties scale with different astrophysical parameters of their parent cluster (Knigge et al., 2009; Leigh et al., 2011; Jadhav & Subramaniam, 2021), which, in principle, could help constrain which of the formation pathways dominates. However, it is important to mention that these formation mechanisms most likely coexist, i.e. they operate simultaneously (Mapelli et al., 2006; Leigh et al., 2011).
@@
Ferraro et al. (2009) found a double sequence of BSSs in the GC M30: a blue and a red one, both roughly aligning with the zero-age MS (ZAMS). The authors argue that the distinction between these two BSS sequences may be due to their origin: collisional BSSs (Col-BSSs), which are deemed more massive, hotter and, therefore, appear bluer in the CMD, whereas MT-BSSs would be less massive, mildly cooler than Col-BSSs and, therefore, appear as a redder BSS sequence. This is not an isolated case, double/multiple sequences of BSSs have been reported for more GCs such as NGC 362 (Dalessandro et al., 2013), NGC 1261 (Simunovic et al., 2014), M 15 (Beccari et al., 2019) and NGC 6256 (Cadelano et al., 2022). The explanation given for all these cases leads to a core-collapse scenario of the host GC in which the central interactions between stars increase, resulting in the formation of BSSs. Therefore, a double/multiple BSS sequence could be a vestige of this event. It also has been found that some BSSs do show the presence of a hot companion (Gosnell et al., 2015; Sindhu et al., 2019), providing supporting evidence for the MT formation scenario in action. Linking these concepts, Dattatrey et al. (2023, hereafter D23) have analyzed the spectral energy distribution (SED) for multiple BSSs in the GC NGC 362. They demonstrate that BSSs that do and do not have a hot companion lie on the red and blue BSS sequence, respectively, in agreement with explanations provided by Dalessandro et al. (2013) for the same GC. More recently, a double BSS sequence has also been detected in the OC Berkeley 17 by Rao et al. (2023a). Given the lack of core collapse in OCs, and, therefore, a considerable attenuation of the Col-BSS formation channel, the authors argue that the presence of this double BSS sequence may be due to a significant difference in BSS rotational velocities, the presence of multiple stellar populations (MSP), or due to extinction effects (for more details see Rao et al., 2023a, and references therein).
1.2 Combining Gaia and HST data
@@
The arrival of the Gaia mission (Gaia Collaboration et al., 2016) with data for almost 1.7 billion objects has opened new opportunities to explore the kinematics of our galaxy and has helped to identify and confirm the cluster membership of stars in GCs (Vasiliev & Baumgardt, 2021) and OCs (Cantat-Gaudin et al., 2018; Dias et al., 2021). However, Gaia faces limitations in overcrowded regions – large errors for its astrometric parameters since it cannot reliably resolve sources – such as GC centres. Nevertheless, BSSs studies in crowded regions are covered by HST data, which can lead to high accuracy measurements in the densest regions (e.g. Baldwin et al., 2016) and, also, since HST features a broader SED coverage compared with Gaia, BSSs can be easily identified through the so-called “UV-route” (Ferraro et al., 1997; Raso et al., 2017; Ferraro et al., 2018, and references therein).
@@
It is particularly interesting to analyze the radial distribution of BSSs, especially in the inner regions of clusters. Since BSSs are more massive compared with other stars in the same cluster, they tend to undergo efficient mass segregation, i.e. “sink” to the inner parts of the cluster (Contreras Ramos et al., 2012). It, is therefore, expected that these peculiar stars can be used to measure the dynamical state of a cluster, i.e., they can be used as “dynamical clocks” in GCs (Ferraro et al., 2012, 2023) and OCs (Rao et al., 2021, 2023b).
@@
Gaia Data Release 3 (Gaia Collaboration et al., 2023, hereafter DR3) for OCs does not suffer from overcrowding, even in the core of OCs. For example, using Gaia Data Release 2 (Gaia Collaboration et al., 2018, hereafter DR2), Rain et al. (2021, hereafter R21) created a catalog for 408 BSSs in different OCs, while Jadhav & Subramaniam (2021, hereafter JS21) have studied the scaling relations of BSSs and their parent OCs. Gaia DR2 has also opened new opportunities to study GCs and their kinematics (e.g., Vasiliev & Baumgardt, 2021). In the past, several HST-based studies have characterized BSSs in GCs (Simunovic & Puzia, 2016). However, the limitation of HST-based studies was that it can only cover the inner regions of clusters, i.e. from the center up to about the half-light radius, . For example, the “HST UV Globular Cluster Survey” (Piotto et al., 2015; Nardiello et al., 2018, also known as “HUGS”) data covers this region and does not include the outskirts of the target GCs. Therefore, combining Gaia DR3 with HST will help us to analyze the radial distribution of BSSs in GCs from the core to outer radii and to determine whether the GCs are dynamically settled or not. In the case of OCs, Gaia DR3 alone can be used to study their dynamical evolution. Hence, a homogeneous catalog of BSSs in GCs and OCs is necessary to understand their formation channels and to examine the dynamical evolution of their host clusters.
@@
In this work, we use Gaia DR3 to create a new homogeneous BSSs catalogue of GCs and OCs, in contrast to previous ones (Rain et al., 2021; Simunovic & Puzia, 2016). This work includes BSSs that were previously not covered due to the radial coverage limitations of prior work. This study also provides various newly derived astrophysical parameters of BSSs from color–temperature relation, isochrone models, and DR3 spectra.
@@
This paper is organized as follows. In Section 2, we describe the selection of cluster members and the BSSs selection. We introduce the different methods applied to obtain different astrophysical parameters for our identified BSS cluster members. In Section 3, we present our results for the total number of identified BSSs and their fraction with respect to their parent cluster MS stars, as well as the distributions for the derived astrophysical BSS parameters. We analyze the results and compare them with previous ones in the literature, and discuss the implications. We finally discuss possible uses and extensions for the identified BSS sample. In Section 4, we summarize the main results of this work.
@@ 2 Data Selection
To select the clusters that are studied in this work, we first select GCs and OCs that have a significant amount of members based on previous studies. We select all the GCs from Vasiliev & Baumgardt (2021) catalog with at least 1000 identified member stars and all the OCs from Dias et al. (2021) catalog with at least identified 350 member stars. We find 41 GCs and 129 OCs following the above criteria.
2.1 Stars Cluster Membership
To identify cluster members we apply a technique developed and used by Cordoni et al. (2018, hereafter C18), with a minor change for OCs (see step 2.), which has also been used by Cordoni et al. (2020) for GCs. Our procedure is summarized as follows:
@@ 2.1 Stars Cluster Membership
First, we select data from Gaia DR3 based on coordinates provided by Vasiliev & Baumgardt (2021) for GCs and Dias et al. (2021) for OCs. As the data query radius we use the Rscale value –the scale radius of Gaia-detected cluster members–, plus , from Vasiliev & Baumgardt (2021) for GCs. For OCs we select stars inside –the radius containing half of the identified members from Dias et al. (2021)–, with a maximum apparent radius of .
@@
We only keep data with G magnitudes in the range between and , with an error in proper motion (in both components along RA and Dec) lower than and the renormalised unit weight error (ruwe) lower than 1.4.
@@
We plot the G magnitude vs. and split the magnitude component into bins with bin size of mag. In every bin, we compute the median values for mag and the parameter, and the rms () for . We then create points with coordinates G median (vertical axis) and (horizontal axis). We interpolate all the points created in this way and we keep all the values within under this interpolation.
@@
We repeat the same procedure as the previous step, but with the parallax () instead of . For every bin, we create two boundaries with coordinates points based median mag (vertical axis) and median (horizontal axis), interpolate them and keep data between these boundaries.
We iterate this procedure a total of 3 times starting from step 3. For a more detailed description of this procedure, we encourage the reader to view section 2 and figure 1 from C18 and Cordoni et al. (2020). Figure 1 shows the distribution of the member stars in our GC and OC sample. We have also compared our cluster members with the literature (Dias et al., 2021; Vasiliev & Baumgardt, 2021; Hunt & Reffert, 2023; Cantat-Gaudin et al., 2020a) and found a very good match with them. The small discrepancy in the crossmatch is found mostly toward the fainter magnitudes (G18). The details of this crossmatch are described in the Appendix A.
2.2 Selection of Blue Straggler Stars
We create a CMD ridge-line for each cluster using the median color and magnitude in bins in the vs. CMD and select MS, sub-giant branch (SGB) and red giant branch (RGB) stars. We prefer over the magnitude in the CMD since it is expected that BSSs will appear brighter at bluer wavelengths across the Gaia SED coverage (see, e.g., Raso et al., 2017). However, the difference of using or does not change our final sample size as covers approximately –, while covers – (Rodrigo et al., 2012; Rodrigo & Solano, 2020), and given the lack of significant near-IR emission of BSS star the wider coverage of the filter mainly adds noise and not useful flux information for BSSs (see e.g. D23, ).
@@ 2.2 Selection of Blue Straggler Stars
Finally, we recreate a zero-age main sequence (ZAMS) keeping the stellar population parameters for the best-fitting isochrone and shifting it to bluer colors, i.e. twice the quantity we have shifted the best-fitting isochrone before. The only 2 differences in the BSS selection between GCs and OCs are in the color selection: i) for GCs we have set a cut in the color up to 0.5 mag to colors bluer than the MSTO to avoid selecting evolved horizontal branch (HB) stars (see, e.g., Culpan et al., 2021). We notice that with this criteria, the selected BSS candidates are found to be located at least five sigma away from the BHB ridge line for all GCs. However, we do not apply this criterion to OCs, as they do have any BHB populations. ii) For OCs, we have applied a cut in color to mag to avoid selecting RGB stars for younger clusters.
@@
We detect 4399 BSSs in our cluster sample. All GCs show the presence of BSSs, with a total of 3965 detected BSSs ( of the sample). On the other hand, only 42 out of 129 selected OCs in our sample show the presence of BSSs, with a total of 434 BSSs ( of the sample). Comparing our BSS members with the literature, we find that more than 90% of all BSSs are also classified as members in previous studies.

### v2

@@ Abstract
Context. The study of blue straggler stars (BSSs) provides insight into the mechanisms of stellar mass exchange during binary stellar evolution and the complex gravitational interactions within dense stellar systems. In combination, they enhance our understanding of the possible life cycles of stars and the evolutionary pathways of star clusters.
Aims. We study the populations of BSSs in 41 globular clusters (GCs) and 42 open clusters (OCs) based on photometry, proper motion, and parallax from the Gaia Data Release 3 (DR3). We confirm their cluster membership. We find a total of 4399 BSSs: All GCs show BSSs (3965 or of the sample), whereas only 42 out of 129 studied OCs show BSSs (434 or of the sample). Clusters younger than Myr do not host any BSSs.
Methods. We derived their astrophysical parameters such as effective temperature, surface gravity, and mass based on color-temperature relations, isochrone models, and Gaia DR3 spectroscopy (if available). We found values for and and an average mass of and for GCs and OCs, respectively. We finally computed the difference of the BSS mass and the main-sequence turn-off (MSTO) mass of its respective cluster, normalized by the MSTO mass, for every identified BSS.
Results. Based on this parameter and on the BSSs ages derived from isochrone models, we find that i) GC BSSs that are most likely to be formed through collisions show a boost for ages Gyr. This agrees with the ages for core-collapse events in GCs that were reported in previous studies. We also find ii) a double sequence for GC BSSs that might indicate a pre- or post-merger or close-binary scenario.
@@
blue stragglers – stars: fundamental parameters — globular clusters: general – open clusters and associations: general1 Introduction
@@
Blue straggler stars (BSSs) are stars that appear to be younger and more massive than the bulk of other stars in their host stellar population. They are characterized by their location in the color-magnitude diagram (CMD), where they appear hotter and bluer than their main-sequence (MS) neighbors.
@@
Blue straggler stars were first identified by Sandage (1953) through his observations of the globular cluster (GC) Messier 3 (M3), and some years later, they were reported in the open cluster (OC) NGC 7786 by Burbidge & Sandage (1958). Ever since, their presence has been noted in distinct structures: as field stars (Preston & Sneden, 2000; Jofré et al., 2016), in OB associations (Mathys, 1987), in the Galactic bulge (Clarkson et al., 2011), and in dwarf galaxies (Da Costa, 1984; Mapelli et al., 2009).
@@
Blue straggler stars challenge the standard theory of stellar evolution, which states that the more massive a star, the faster it evolves. For this reason, these stars should have evolved off the MS branch a long time ago, but they still populate this branch as an apparent extension of the MS in a CMD. Three scenarios have been proposed to explain their formation: i) Mass transfer (MT) in binary systems (McCrea, 1964), where a star receives material from its companion. This extends its life. ii) Stars formed through collisions or mergers (Hills & Day, 1976; Lombardi et al., 2002). iii) In hierarchical triple systems, magnetized winds or stellar evolution cause the loss of angular momentum in a multiple system, which leads to collisions (or MT) between the stars that conform the system (Perets & Fabrycky, 2009; Naoz & Fabrycky, 2014). It has been shown that the BSS properties scale with different astrophysical parameters of their parent cluster (Knigge et al., 2009; Leigh et al., 2011; Jadhav & Subramaniam, 2021), which might help us in principle to constrain which of the formation pathways dominates. It is important to mention, however, that these formation mechanisms most likely coexist, that is, they operate simultaneously (Mapelli et al., 2006; Leigh et al., 2011).
@@
Ferraro et al. (2009) found a double sequence of BSSs in the GC M30: a blue and a red BSS sequence, which were both roughly aligned with the zero-age MS (ZAMS). The authors argued that the distinction between these two BSS sequences may be due to their origin: collisional BSSs (Col-BSSs), which are deemed more massive and hotter and therefore appear bluer in the CMD, whereas MT-BSSs would be less massive, mildly cooler than Col-BSSs, and would therefore appear as a redder BSS sequence. This is not an isolated case. Double/multiple sequences of BSSs have been reported for more GCs, such as NGC 362 (Dalessandro et al., 2013), NGC 1261 (Simunovic et al., 2014), M 15 (Beccari et al., 2019), and NGC 6256 (Cadelano et al., 2022). The explanation given for all these cases leads to a core-collapse scenario of the host GC in which the central interactions between stars increase, resulting in the formation of BSSs. Therefore, a double/multiple BSS sequence could be a vestige of this event. It also has been found that some BSSs show a hot companion (Gosnell et al., 2015; Sindhu et al., 2019), which provides supporting evidence for the MT formation scenario in action. Linking these concepts, Dattatrey et al. (2023, hereafter D23) have analyzed the spectral energy distribution (SED) for multiple BSSs in the GC NGC 362. They demonstrated that BSSs with and without a hot companion lie on the red and blue BSS sequence, respectively, which agrees with the explanations provided by Dalessandro et al. (2013) for the same GC. More recently, a double BSS sequence has also been detected in the OC Berkeley 17 by Rao et al. (2023a). Given the lack of core collapse in OCs and the accordingly considerable attenuation of the Col-BSS formation channel, the authors argued that this double BSS sequence might be due to a significant difference in BSS rotational velocities, the presence of multiple stellar populations (MSP), or to extinction effects (for more details see Rao et al., 2023a, and references therein).
@@
The arrival of the Gaia mission (Gaia Collaboration et al., 2016) with data for almost 1.7 billion objects has opened new opportunities to explore the kinematics of our Galaxy and has helped us to identify and confirm the cluster membership of stars in GCs (Vasiliev & Baumgardt, 2021) and OCs (Cantat-Gaudin et al., 2018; Dias et al., 2021). Gaia faces limitations in overcrowded regions (large errors for its astrometric parameters because it cannot reliably resolve sources), however, such as GC centers. Nevertheless, BSS studies in crowded regions are covered by HST111Hubble Space Telescope data, which can lead to high-accuracy measurements in the densest regions (e.g. Baldwin et al., 2016), and because HST features a broader SED coverage than Gaia, BSSs can also be easily identified through the so-called UV-route (Ferraro et al., 1997; Raso et al., 2017; Ferraro et al., 2018, and references therein).
@@
It is particularly interesting to analyze the radial distribution of BSSs, especially in the inner regions of clusters. Since BSSs are more massive compared with other stars in the same cluster, they tend to undergo efficient mass segregation, that is, they sink to the inner parts of the cluster (Contreras Ramos et al., 2012). It is therefore expected that these peculiar stars can be used to measure the dynamical state of a cluster, that is, they can be used as dynamical clocks in GCs (Ferraro et al., 2012, 2023) and OCs (Rao et al., 2021, 2023b).
@@
Gaia Data Release 3 (Gaia Collaboration et al., 2023, hereafter DR3) for OCs does not suffer from overcrowding, even in the core of OCs. For example, using Gaia Data Release 2 (Gaia Collaboration et al., 2018, hereafter DR2), Rain et al. (2021, hereafter R21) created a catalog for 408 BSSs in different OCs, while Jadhav & Subramaniam (2021, hereafter JS21) have studied the scaling relations of BSSs and their parent OCs. Gaia DR2 has also opened new opportunities for studying GCs and their kinematics (e.g., Vasiliev & Baumgardt, 2021). Several HST-based studies have previously characterized BSSs in GCs (Simunovic & Puzia, 2016). The limitation of HST-based studies was, however, that they can only cover the inner regions of clusters, that is, from the center up to about the half-light radius, . For example, data from the HST UV Globular Cluster Survey (Piotto et al., 2015; Nardiello et al., 2018, also known as “HUGS”) cover this region and do not include the outskirts of the target GCs. Therefore, combining Gaia DR3 with HST will help us to analyze the radial distribution of BSSs in GCs from the core to the outer radii and to determine whether the GCs are dynamically settled. In the case of OCs, Gaia DR3 alone can be used to study their dynamical evolution. Hence, a homogeneous catalog of BSSs in GCs and OCs is necessary to understand their formation channels and to examine the dynamical evolution of their host clusters.
@@
We use Gaia DR3 to create a new homogeneous BSSs catalog of GCs and OCs that differs from previous catalogs (Rain et al., 2021; Simunovic & Puzia, 2016). This work includes BSSs that were previously not covered due to the radial coverage limitations of prior work. This study also provides various newly derived astrophysical parameters of BSSs from the color–temperature relation, isochrone models, and DR3 spectra.
@@
This paper is organized as follows. In Section 2 we describe the selection of cluster members and the selection of BSSs. We introduce the different methods we applied to obtain different astrophysical parameters for the BSS cluster members we identified. In Section 3 we present our results for the total number of identified BSSs and their fraction with respect to their parent cluster MS stars, as well as the distributions for the derived astrophysical BSS parameters. We analyze the results and compare them with previous results in the literature, and we discuss the implications. We finally discuss possible uses and extensions for the identified BSS sample. In Section 4 we summarize the main results of this work.
@@ 2 Data selection
To select the clusters studied in this work, we first select GCs and OCs that have a significant number of members based on previous studies. We select all GCs from the catalog by Vasiliev & Baumgardt (2021) with at least 1000 identified member stars and all OCs from the catalog by Dias et al. (2021) with at least 350 identified member stars. We find 41 GCs and 129 OCs following the above criteria.
2.1 Star cluster membership
To identify cluster members, we apply a technique developed and used by Cordoni et al. (2018, hereafter C18), with a minor change for OCs (see step 2.), which has also been used by Cordoni et al. (2020) for GCs. Our procedure is summarized as follows:
@@ 2.1 Star cluster membership
First, we select data from Gaia DR3 based on coordinates provided by Vasiliev & Baumgardt (2021) for GCs and Dias et al. (2021) for OCs. As the data query radius, we use the Rscale value –the scale radius of Gaia-detected cluster members–, plus , from Vasiliev & Baumgardt (2021) for GCs. For OCs we select stars inside –the radius containing half of the identified members from Dias et al. (2021)–, with a maximum apparent radius of .
@@
We only keep data with G magnitudes in the range between and , with an error in proper motion (in both components along RA and Dec) lower than and the renormalized unit weight error (ruwe) lower than 1.4.
@@
We plot the G magnitude versus and split the magnitude component into bins with bin size of mag. In every bin, we compute the median values for mag and the parameter, and the rms () for . We then create points with coordinates G median (vertical axis) and (horizontal axis). We interpolate all the points created in this way and we keep all the values within under this interpolation.
@@
We repeat the same procedure as the previous step, but with the parallax () instead of . For every bin, we create two boundaries with coordinate points-based median mag (vertical axis) and median (horizontal axis), interpolate them, and keep the data between these boundaries.
We iterate this procedure a total of 3 times starting from step 3. For a more detailed description of this procedure, we encourage the reader to view section 2 and figure 1 from C18 and Cordoni et al. (2020). Figure 1 shows the distribution of the member stars in our GC and OC sample. We also compared our cluster members with the literature (Dias et al., 2021; Vasiliev & Baumgardt, 2021; Hunt & Reffert, 2023; Cantat-Gaudin et al., 2020a) and found a very good match with them. The small discrepancy in the crossmatch is found mostly towards the fainter magnitudes (G18). The details of this crossmatch are described in Appendix A.
2.2 Selection of BSSs
We created a CMD ridge-line for each cluster using the median color and magnitude in bins in the versus CMD and select MS, subgiant branch (SGB) and red giant branch (RGB) stars. We prefer over the magnitude in the CMD since it is expected that BSSs will appear brighter at bluer wavelengths across the Gaia SED coverage (see, e.g., Raso et al., 2017). The difference of using or does not change our final sample size, however, because covers approximately –, while covers – (Rodrigo et al., 2012; Rodrigo & Solano, 2020), and given the lack of significant near-IR emission of BSS star the wider coverage of the filter mainly adds noise and not useful flux information for BSSs (see e.g. D23, ).
@@ 2.2 Selection of BSSs
Finally, we recreate a zero-age main sequence (ZAMS) keeping the stellar population parameters for the best-fitting isochrone and shifting it to bluer colors, i.e. twice the quantity we shifted the best-fitting isochrone before. The only 2 differences in the BSS selection between GCs and OCs are in the color selection: i) for GCs we set a cut in the color up to 0.5 mag to colors bluer than the MSTO to avoid selecting evolved horizontal branch (HB) stars (see, e.g., Culpan et al., 2021). We notice that with this criteria, the selected BSS candidates are found to be located at least five sigma away from the BHB ridge line for all GCs. We did not apply this criterion to OCs, however, because they have no BHB populations. ii) For OCs, we applied a cut in color to mag to avoid selecting RGB stars for younger clusters.
@@
We detect 4399 BSSs in our cluster sample. All GCs show BSSs, with a total of 3965 detected BSSs ( of the sample). On the other hand, only 42 out of 129 selected OCs in our sample show BSSs, with a total of 434 BSSs ( of the sample). Comparing our BSS members with the literature, we find that more than 90% of all BSSs are also classified as members in previous studies.